Use a double integral to find the area enclosed by one loop of the four-leaved rose r =cos(2 theta).
1
Expert's answer
2020-10-28T17:43:56-0400

Consider the four leaved rose "r=cos(2\theta)"


For "r=0" , solve "\theta" as,


"cos(2\theta)=0"


"2\theta=\frac{\pi}{2}" in "[0,\frac{\pi}{2}]"


"\theta=\frac{\pi}{4}"


The sketch of the curve is as shown in the figure below:





The area enclosed by one loop of the rose is evaluated as,


Area"(A)=2\int_{\theta=0}^{\frac{\pi}{4}}\int_{r=0}^{cos(2\theta)}rdrd\theta" ....[using symmetry]


"=2\int_{\theta=0}^{\frac{\pi}{4}}[\frac{r^2}{2}]_{r=0}^{cos(2\theta)}d\theta"


"=2(\frac{1}{2})\int_{\theta=0}^{\frac{\pi}{4}}cos^2(2\theta)d\theta"


"=\int_{\theta=0}^{\frac{\pi}{4}}[\frac{1+cos(4\theta)}{2}]d\theta"


"=\frac{1}{2}\int_{\theta=0}^{\frac{\pi}{4}}(1+cos(4\theta))d\theta"


"=\frac{1}{2}[\theta+\frac{sin(4\theta)}{4}]_{0}^{\frac{\pi}{4}}"


"=\frac{1}{2}[\frac{\pi}{4}+\frac{0-0}{4}]"


"=\frac{1}{2}[\frac{\pi}{4}+0]"


"=\frac{\pi}{8}"


Therefore, the area of one loop of the rose is Area"(A)=\frac{\pi}{8}" square units.